OpenStudy (ashley7777):

2sinx/3+sqr3=0 solve for x

OpenStudy (studygurl14):

Hello @ashley7777 ! Welcome to Openstudy! Is it like this? \(\Large \frac{2\sin x}{3}+\large \sqrt{3}=0\)

OpenStudy (ashley7777):